# Daniele Bartoli1, Mohit Pal2, Pantelimon Stănică3, Tommaso Toccotelli1

### v1

@@ 2 Main result
It is easy to observe that when $m$ is even then $C_{u}$ is not a permutation as in this case $C_{u}(X,0,0)=(X^{3},0,0)$ and the function $X\mapsto X^{3}$ is $3$-to-$1$. In what follows we assume that $m$ is odd. Note that $C_{u}$ is a permutation if and only if the equation
$C_{u}(X+\alpha,Y+\beta,Z+\gamma)+C_{u}(X,Y,Z)=0$
has only the trivial solutions $\{(x,y,z,0,0,0)\ :\ x,y,z\in\mathbb{F}_{2^{m}}\}$.
Such a condition reads
@@
It is straightforward to see that the first equation of the above system has no solution $(\overline{x},\overline{y},\overline{z})\in\mathbb{F}_{2^{m}}^{3}$ as $m$ is odd.
@@
which can be further simplified by replacing $X\mapsto\alpha X$ and $Y\mapsto\beta Y$, i.e., System (3) is equivalent to the following system
@@
Now, squaring the first equation and putting $X^{2}=Y+1$ into it, we have
@@
Thus, System (4) is equivalent to the following system
@@
Let
@@
Proposition 4.
Each element of
$\Theta:=$$\{(x,y,z,a,b,c)\in\mathbb{F}_{q}^{6}:(b+y)h(b,x,y,z)\neq 0,\ c=f(a,b,x,y,z),$
$\qquad\qquad\qquad\qquad a=g(b,x,y,z)/h(b,x,y,z),\sum_{i=0}^{21}a_{i}(b,y,z)x^{i}=0\}$
is a solution of System (2).
Proof.
This is easily checked via direct computations in MAGMA [6]. In particular, from the first equation in System (2) one gets $\gamma=f(\alpha,\beta,X,Y,Z)$, and combining this condition with the second equation one can check that $\alpha=g(\beta,X,Y,Z)/h(\beta,X,Y,Z)$ holds. Finally, let $H(\beta,X,Y,Z)$ be the polynomial obtained by combining the third equation in System (2) with $\gamma=f(\alpha,\beta,X,Y,Z)$ and $\alpha=g(\beta,X,Y,Z)/h(\beta,X,Y,Z)$, after clearing the denominators. It can be checked that the polynomial $\sum_{i=0}^{21}a_{i}(\beta,Y,Z)X^{i}$ is a factor of $H(\beta,X,Y,Z)$, thus completing the proof. ∎
Proposition 5.
Let $\mathcal{V}\subset\mathbb{A}^{6}(\mathbb{F}_{q})$ be the variety defined by
$\begin{cases}X_{3}=f(X_{1},X_{2},X_{4},X_{5},X_{6})\\ X_{1}=g(X_{2},X_{4},X_{5},X_{6})/h(X_{2},X_{4},X_{5},X_{6})\\ \sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}=0.\end{cases}$
Then $\mathcal{V}$ contains an absolutely irreducible component defined over $\mathbb{F}_{q}$, distinct from $X_{1}=X_{2}=X_{3}=0$, of degree at most $27$, and not contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$.
Proof.
Since $\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$ is a homogeneous polynomial of degree $24$ it defines a surface $\mathcal{S}$ in $\mathbb{P}^{3}(\mathbb{F}_{q})$. First note that $X_{2}$ is not a factor of $\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$, otherwise $u\in\mathbb{F}_{4}^{*}$, a contradiction. Also, $\mathcal{S}$ contains an absolutely irreducible non-repeated component defined over $\mathbb{F}_{q}$, since its intersection with the plane $X_{4}=0$ is $a_{0}(X_{2},X_{5},X_{6})=0$ and
$X_{2}\mid a_{0}(X_{2},X_{5},X_{6})\textrm{ and }X_{2}^{2}\nmid a_{0}(X_{2},X_{5},X_{6}),$
by Lemma 2 there exists an absolutely irreducible component $\mathcal{S}^{\prime}$ defined over $\mathbb{F}_{q}$ whose intersection with $X_{4}=0$ is $X_{2}=0$.
Since the intersection between $X_{4}=0$ and $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ is
$u^{5}X_{6}^{2}(X_{2}+X_{5})^{3}((u^{2}+u^{5})X_{2}^{3}X_{5}^{4}+(u^{2}+u^{5})X_{2}^{2}X_{5}^{5}+(u^{2}+u^{5})X_{2}X_{5}^{6}+u^{5}X_{5}^{7}+X_{6}^{7})=0,$
$\mathcal{S}^{\prime}$ cannot be contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$.
Clearly, $\mathcal{S}^{\prime}$ extends to a variety $\mathcal{V}^{\prime}$ contained in $\mathcal{V}$, absolutely irreducible and $\mathbb{F}_{q}$-rational (by considering the two extra equations $X_{3}=f(X_{1},X_{2},X_{4},X_{5},X_{6})$ and $X_{1}=g(X_{2},X_{4},X_{5},X_{6})/h(X_{2},X_{4},X_{5},X_{6}))$. Since $\mathcal{S}^{\prime}$ is different from $X_{2}=0$, $\mathcal{V}^{\prime}$ is different from $X_{1}=X_{2}=X_{3}=0$. The degree of $\mathcal{V}^{\prime}$ is upper bounded by the degree of $\mathcal{V}$, that is at most $273^{3}=$ 27. ∎
The following theorem is the main result of our paper.
Theorem 6.
If $m\geq 25$, then Conjecture 3 is true.
Proof.
By Proposition 5 and Theorem 1, if $m$ is large enough, the absolutely irreducible $\mathbb{F}_{q}$-rational component $\mathcal{V}^{\prime}$ (and thus $\mathcal{V}$) possesses roughly $q^{3}$ of $\mathbb{F}_{q}$-rational points. Since $\mathcal{V}^{\prime}$ is not $X_{1}=X_{2}=X_{3}=0$ and it not contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$, the set $\Theta$ is not empty and the validity of the conjecture follows.
The bound $m\geq 25$ can be shown via Theorem 1. First, notice that points of $\mathcal{V}^{\prime}$ not belonging to $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ correspond to quadruples $(x_{2},x_{4},x_{5},x_{6})\in\mathbb{F}_{q}^{4}$ such that $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $(x_{2}+x_{5})h(x_{2},x_{4},x_{5},x_{6})\neq 0$. It can be easily verified that the intersections of $\mathcal{S}$ and $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ with $X_{4}=0$ are coprime: this shows that they do not share any component. Also, since the degree of these surfaces is $24$ and $12$ respectively, their intersection is a projective space curve of degree $288$ and it contains at most $288q$ projective $\mathbb{F}_{q}$-rational points; see [11]. This shows the existence of at most $288q(q-1)$ nonzero quadruples $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{F}_{q}^{4}$ such that $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $(x_{2}+x_{5})h(x_{2},x_{4},x_{5},x_{6})=0$. Arguing similarly, there are at most $24q(q-1)$ nonzero quadruples $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{F}_{q}^{4}$ satisfying $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $x_{2}=0$ (and thus lying on $X_{1}=X_{2}=X_{3}=0$). By Theorem 1 the number of $\mathbb{F}_{q}$-rational points in $\Theta$ is at least
$q^{3}-25\cdot 26\cdot q^{5/2}-5\cdot(27)^{13/3}q^{2}-312q(q-1),$
that is larger than 00 when $m\geq 25$ and thus Conjecture 3 holds true. ∎
3 Conclusion
In this paper, using algebraic geometry tools, we show a conjecture of Beierle, Carlet, Leander, Perrin from 2022, when the underlying field has dimension $m\geq 25$.
Acknowledgements
The third-named author (PS) would like to thank the first-named author (DB) for the invitation at the Dipartimento di Matematica e Informatica at Università degli Studi di Perugia, Italy, and the great working conditions while this paper was being written. The first-named author (DB) thanks the Italian National Group for Algebraic and Geometric Structures and their Applications (GNSAGA—INdAM) which supported the research.

### v2

@@ 5 Complete characterization for $m=3$
Note that for even $m$, $C_{u}$ fails to be a permutation, since $C_{u}(X,0,0)=(X^{3},0,0)$ and the function $X\mapsto X^{3}$ is 33-to-11. In what follows we assume that $m$ is odd.
We analyze when $C_{u}$ fails to be a permutation by studying the differential equation
$C_{u}(X+\alpha,Y+\beta,Z+\gamma)+C_{u}(X,Y,Z)=(0,0,0).$
The function $C_{u}$ is a permutation if and only if this equation has only trivial solutions $\{(x,y,z,0,0,0):x,y,z\in\mathbb{F}_{2^{m}}\}$. Expanding using the fact that $(A+B)^{3}=A^{3}+B^{3}$ in characteristic 2, this condition becomes
@@
It is routine to verify that the first equation of the above system has no solution $(\overline{x},\overline{y},\overline{z})\in\mathbb{F}_{2^{m}}^{3}$ as $m$ is odd (it is enough to observe that is equivalent to $X^{2}+X+1=0$ via $X\mapsto\alpha X$, and such equation has solutions in $\mathbb{F}_{4}$).
@@
which can be further simplified by replacing $X\mapsto\alpha X$ and $Y\mapsto\beta Y$, i.e., System (4) is equivalent to:
@@
Next, squaring the first equation and putting $X^{2}=Y+1$ into it, we have
@@
Thus, System (5) is equivalent to:
@@
Let $f(\alpha,\beta,X,Y,Z)$, $g(\beta,X,Y,Z)$, $h(\beta,X,Y,Z)$, $a_{i}(\beta,Y,Z)$, $i=0,\ldots,21$, be defined as in the Appendix.
Proposition 6.1.
Define
$\Theta:=$ $\{(x,y,z,a,b,c)\in\mathbb{F}_{2^{m}}^{6}:(b+y)h(b,x,y,z)\neq 0,\ c=f(a,b,x,y,z),$
$\qquad\qquad\qquad\qquad a=g(b,x,y,z)/h(b,x,y,z),\sum_{i=0}^{21}a_{i}(b,y,z)x^{i}=0\}.$
Each element of $\Theta$ is a solution of System (3). Conversely, every non-trivial solution of System (3) with $\beta\neq 0$ and $(b+y)h(b,x,y,z)\neq 0$ corresponds to an element in $\Theta$.
Proof.
We establish the equivalence using computational algebraic geometry, verified in MAGMA [6]. From the first equation of System (3),
$\alpha X^{2}+\alpha^{2}X+u\beta^{2}Z=\alpha^{3}+u\beta^{2}\gamma,$
assuming $\beta\neq 0$, we solve for $\gamma$, obtaining
$\gamma=f(\alpha,\beta,X,Y,Z):=\frac{\alpha^{3}+\alpha^{2}X+\alpha X^{2}+\beta^{2}Zu}{\beta^{2}u+Y^{2}u}.$
Substituting this expression into the second equation of System (3) and clearing denominators yields a polynomial equation that can be solved for $\alpha$, and get
$\alpha=\frac{g(\beta,X,Y,Z)}{h(\beta,X,Y,Z)},$
where the condition $h(\beta,X,Y,Z)\neq 0$ ensures the solution is well-defined. By the theory of Gröbner bases and elimination ideals, when we eliminate $\alpha$ and $\gamma$ from the three-equation system, the resulting polynomial in $(\beta,X,Y,Z)$ encodes all possible solutions. Substituting both expressions into the third equation and clearing denominators yields a polynomial $H(\beta,X,Y,Z)$ that vanishes on the solution set.
Computational factorization shows
$H(\beta,X,Y,Z)=\left(\sum_{i=0}^{21}a_{i}(\beta,Y,Z)X^{i}\right)\cdot R(\beta,X,Y,Z),$
which can be verified using the code in [11]. Therefore, any zero of $\sum_{i=0}^{21}a_{i}(\beta,Y,Z)X^{i}$ with $(Y+\beta)h(\beta,X,Y,Z)\neq 0$ lifts to a unique solution of System (3) via the formulas for $\alpha$ and $\gamma$. The explicit forms of the polynomials $f,g,h,$ and $a_{i}$ are listed in Appendix A. ∎
To apply projective geometry techniques, we work in projective coordinates $(X_{2}:X_{4}:X_{5}:X_{6})$ corresponding to the affine coordinates $(\beta,X,Y,Z)$ used in Proposition 6.1, with the identification $(X_{2}:X_{4}:X_{5}:X_{6})=(\beta:X:Y:Z)$ in affine charts.
Proposition 6.2.
Let $\mathcal{S}\subset\mathbb{P}^{3}(\overline{\mathbb{F}_{q}})$ be the $\mathbb{F}_{q}$-rational surface defined by
$\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}=0.$
Then $\mathcal{S}$ contains an absolutely irreducible component defined over $\mathbb{F}_{q}$, distinct from $X_{2}=0$, of degree at most $24$, and not contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$.
Proof.
First note that $X_{2}$ is not a factor of $\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$, otherwise $u\in\mathbb{F}_{4}^{*}$, a contradiction. To prove that the surface $\mathcal{S}$ contains a non-repeated, absolutely irreducible component defined over $\mathbb{F}_{q}$, we analyze its defining polynomial $F(X_{2},X_{4},X_{5},X_{6})=\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$.
Consider the intersection of $\mathcal{S}$ with the plane $X_{4}=0$. The resulting curve is defined by $F(X_{2},0,X_{5},X_{6})=a_{0}(X_{2},X_{5},X_{6})$. We observe that
$X_{2}\mid a_{0}(X_{2},X_{5},X_{6})\quad\text{and}\quad X_{2}^{2}\nmid a_{0}(X_{2},X_{5},X_{6}).$
This indicates that $X_{2}$ is a simple (non-repeated) factor of $a_{0}$. Let the factorization of $F$ into absolutely irreducible factors over $\overline{\mathbb{F}_{q}}$ be
$F(X_{2},X_{4},X_{5},X_{6})=H_{1}\cdots H_{r}.$
Setting $X_{4}=0$ yields the factorization $a_{0}=H_{1}|_{X_{4}=0}\cdots H_{r}|_{X_{4}=0}$. Since $X_{2}$ is a simple factor of $a_{0}$, it must divide exactly one factor on the right-hand side, say $H_{k}|_{X_{4}=0}$.
We claim that the component $H_{k}$ is defined over $\mathbb{F}_{q}$. The Frobenius automorphism $\Phi_{q}$ permutes the set of absolutely irreducible factors $\{H_{1},\dots,H_{r}\}$. Since $a_{0}\in\mathbb{F}_{q}[X_{2},X_{5},X_{6}]$, we have $\Phi_{q}(a_{0})=a_{0}$, and since $X_{2}\in\mathbb{F}_{q}$, $\Phi_{q}(X_{2})=X_{2}$. Suppose, for the sake of contradiction, that $H_{k}$ is not defined over $\mathbb{F}_{q}$. Then $\Phi_{q}(H_{k})=H_{j}$ for some index $j\neq k$. Applying $\Phi_{q}$ to the relation $X_{2}\mid H_{k}|_{X_{4}=0}$ yields
$X_{2}=\Phi_{q}(X_{2})\mid\Phi_{q}(H_{k}|_{X_{4}=0})=H_{j}|_{X_{4}=0}.$
This implies that $X_{2}$ divides at least two distinct factors of $a_{0}$, namely $H_{k}|_{X_{4}=0}$ and $H_{j}|_{X_{4}=0}$. Consequently, $X_{2}^{2}$ must divide their product, which divides $a_{0}$. This contradicts the fact that $X_{2}^{2}\nmid a_{0}$. Thus, $\Phi_{q}(H_{k})=H_{k}$, establishing that the absolutely irreducible component $H_{k}$ is defined over $\mathbb{F}_{q}$. Let us denote this component by $\mathcal{S}^{\prime}$.
Finally, we check that $\mathcal{S}^{\prime}$ is not contained in the degenerate locus given by $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$. The intersection of this locus with $X_{4}=0$ is
$u^{5}X_{6}^{2}(X_{2}+X_{5})^{3}((u^{2}+u^{5})X_{2}^{3}X_{5}^{4}+(u^{2}+u^{5})X_{2}^{2}X_{5}^{5}+(u^{2}+u^{5})X_{2}X_{5}^{6}+u^{5}X_{5}^{7}+X_{6}^{7})=0.$
Since $X_{2}$ is not a factor of this polynomial (as the term $X_{6}^{7}$ prevents it), the component $\mathcal{S}^{\prime}$ (which intersects $X_{4}=0$ at $X_{2}=0$) cannot be contained in $(X_{2}+X_{5})h=0$. ∎
Theorem 6.3.
Assume throughout this statement that all parameters lie in $\mathbb{F}_{2^{m}}$ under the standing hypotheses from the preliminaries. If $m\geq 23$ is odd and $u\in\mathbb{F}_{2^{m}}^{*}$, then the function $C_{u}:\mathbb{F}_{2^{m}}^{3}\to\mathbb{F}_{2^{m}}^{3}$ given by
$(X,Y,Z)\mapsto(X^{3}+uY^{2}Z,Y^{3}+uXZ^{2},Z^{3}+uX^{2}Y)$
is not a permutation. This resolves the permutation part of the conjecture posed by Beierle, Carlet, Leander, and Perrin [3].
Proof.
For $C_{u}$ to be a permutation, the system of equations in (3) must only have trivial solutions. We show that for $m\geq 23$, there always exists a non-trivial solution $(\alpha,\beta,\gamma)\neq(0,0,0)$ and a corresponding $(X,Y,Z)$ that satisfies the system. This existence is guaranteed by finding an $\mathbb{F}_{2^{m}}$-rational point on a specific algebraic variety derived from the system.
We analyze the system by algebraic elimination. A solution exists if there is a point $(X,Y,Z,\alpha,\beta,\gamma)$ satisfying the system for some non-zero $(\alpha,\beta,\gamma)$. This problem is equivalent to finding a rational point in the variety defined by the equations in Proposition 6.1.
We use the comments and reductions from our prior propositions. By Proposition 6.2, the surface $\mathcal{S}$ contains an absolutely irreducible $\mathbb{F}_{q}$-rational component, which we denote by $\mathcal{S}^{\prime}$. Our goal is to demonstrate that for a sufficiently large $m$, the size of $\Theta$ is positive, proving the conjecture of Beierle et al. [3].
An element in $\Theta$ is a six-tuple $(x_{1},\dots,x_{6})\in\mathbb{F}_{q}^{6}$ derived from a point on $\mathcal{S}^{\prime}$ that satisfies the system’s equations
$x_{3}=f(x_{1},x_{2},x_{4},x_{5},x_{6}),\qquad h(x_{2},x_{4},x_{5},x_{6})x_{1}=g(x_{2},x_{4},x_{5},x_{6}).$
A solution is considered valid only if it avoids degeneracy, requiring $(x_{2}+x_{5})h(x_{2},x_{4},x_{5},x_{6})\neq 0$. This condition ensures that $x_{1}$ is uniquely determined and non-trivial.
Our strategy is to apply the Lang-Weil estimate from Theorem 2.1 to find a lower bound on the number of $\mathbb{F}_{q}$-rational points on $\mathcal{S}^{\prime}$, and then subtract an upper bound for the number of points that lead to degenerate solutions. These degenerate points lie on the intersection of $\mathcal{S}^{\prime}$ with the surfaces defined by $(X_{2}+X_{5})=0$ or $h(X_{2},X_{4},X_{5},X_{6})=0$.
The degree of $\mathcal{S}^{\prime}$ is at most 24, while the degree is 12 of the surface defined by $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$. Since $\mathcal{S}^{\prime}$ is not a component of this surface (by checking that $\mathcal{S}$ and $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ do not share any component since their intersection with $X_{4}=0$ are coprime), their intersection is a curve of degree at most $24\times 12=288$. By the results in [9], this intersection contains at most $288q$ projective $\mathbb{F}_{q}$-rational points. with the same substitutions and exponent arithmetic, the intersection of $\mathcal{S}^{\prime}$ with the plane $X_{2}=0$ is a curve of degree at most 24, containing at most $24q$ projective points.
In total, the number of ‘degenerate’ projective points (corresponding to degenerate solutions) is at most $(288+24)q=312q$. Each of these corresponds to (q1)(q-1) affine quadruples, so we exclude at most 312q(q1)312q(q-1) quadruples.
From Theorem 2.1, a lower bound on the number of $\mathbb{F}_{q}$-rational points on $\mathcal{S}^{\prime}$ is given by $q^{2}-23\cdot 22\cdot q^{3/2}-5\cdot(24)^{13/3}q$. We first verify that the hypotheses of Theorem 2.1 are satisfied. The theorem requires $q>2(r+1)\delta^{2}$, where $r$ is the dimension and $\delta$ is the degree of the variety.
For our surface $\mathcal{S}^{\prime}$, we have $r=2$ (since $\mathcal{S}^{\prime}$ is a surface in $\mathbb{P}^{3}$) and $\delta\leq 24$ (from Proposition 6.2). Thus the condition becomes
$q>2(r+1)\delta^{2}=2(2+1)(24)^{2}=2\cdot 3\cdot 576=3456.$
Since we assume $m\geq 23$ and $q=2^{m}$, we have $q=2^{23}=8{,}388{,}608\gg 3456$. Therefore, Theorem 2.1 applies, and we obtain the estimate
$|\mathcal{S}^{\prime}(\mathbb{F}_{q})|\geq q^{2}-(\delta-1)(\delta-2)q^{3/2}-5\delta^{13/3}q=q^{2}-23\cdot 22\cdot q^{3/2}-5\cdot(24)^{13/3}q.$
Note that $23\cdot 22=506$ and $24^{13/3}=24^{4}\cdot 24^{1/3}\approx 331{,}776\cdot 2.88\approx 955{,}917$, so the bound is well-defined and valid for $m\geq 23$. The number of corresponding non-zero affine quadruples is (q1)(q-1) times this value.
Thus, a lower bound for $|\Theta|$, the number of valid solutions, is
$t(q)=\left(q^{2}-506q^{3/2}-5\cdot(24)^{13/3}q\right)(q-1)-312q(q-1).$
To ensure $\Theta$ is non-empty, we require $t(q)>0$. Dividing by q(q1)q(q-1), we analyze the positivity of
$s(q):=\frac{t(q)}{q(q-1)}=q-506q^{1/2}-5\cdot(24)^{13/3}-312.$
This inequality holds if $q^{1/2}$ is sufficiently large. The condition $s(x^{2})>0$ holds for $x>2456$, which implies that $|\Theta|>0$ for $q>(2456)^{2}$. This threshold on $q$ is met when $m\geq 23$. Therefore, for $m\geq 23$, the set of valid solutions $\Theta$ is non-empty, and our claim is shown. ∎
Remark 6.4.
The threshold $m\geq 23$ is sufficient for our proof technique but may not be optimal. Computational evidence suggests that $C_{u}$ fails to be a permutation for much smaller values of $m$. The specific threshold arises from the Lang-Weil bound requirement $q>2(r+1)\delta^{2}=3456$ (satisfied by $q=2^{23}=8{,}388{,}608$), combined with our degree estimates for the surface $\mathcal{S}^{\prime}$ (degree $\leq 24$) and intersection bounds with degenerate loci (degree $\leq 288$). Improved degree estimates or alternative geometric approaches might lower this threshold significantly.
7 Conclusion and open problems
We have developed systematic algebraic geometry techniques for establishing non-permutation results for trivariate functions over finite fields. Our main contributions include: first, necessary polynomial conditions (Theorems 4.14.3) that force trivariate functions to be non-permutations, based on differential equation analysis and elimination theory; second, a complete classification showing that for $m=3$, exactly 20 parameter choices yield permutations (Proposition 5.1), falling into two structural families (Type I and Type II) with distinct algebraic properties; and third, the resolution of a conjecture by Beierle et al. (Theorem 6.3) using Lang-Weil bounds and component analysis, proving $C_{u}$ is not a permutation for all odd $m\geq 23$.
These results demonstrate the power of number theory and algebraic geometry methods–particularly elimination theory, projective varieties, and point-counting techniques–for analyzing the permutation properties of multivariate cryptographic functions. The techniques developed here provide a template for studying other classes of vectorial Boolean functions.
Several fundamental questions remain:
Open Problem 7.1.
Do any Type I permutations $(0,a_{2},a_{3},0)$ extend to infinite families for infinitely many $m$?
The 13 Type I permutations found for $m=3$ may be sporadic or may represent infinite families with different structural properties. Our proof techniques do not directly apply to Type I due to vanishing of $a_{1}$ and $a_{4}$.
Open Problem 7.2.
Are there parameter families beyond Types I and II that yield permutations for infinitely many $m$?
Theorem 4.3 shows that generic parameters fail for large $m$, but there may be other special families with hidden algebraic structure that we have not yet identified.
Open Problem 7.3.
Determine whether $C_{u}$ is a permutation for odd $m<23$. Our theoretical approach requires $m\geq 23$ due to the Lang-Weil bound hypothesis $q>2(r+1)\delta^{2}=3456$. The function may behave differently for small $m$, particularly $m=3$. Improved degree estimates or alternative geometric techniques might lower the threshold, or computational methods might resolve specific small cases.
Our current proof establishes $M_{0}=23$. Computational evidence suggests the threshold may be much smaller. Improvements could come from tighter degree bounds, refined Lang-Weil estimates, or direct construction of explicit solutions for specific small $m$.
This work demonstrates the power of algebraic geometry for analyzing multivariate functions. The techniques—particularly elimination theory, projective varieties, and point-counting bounds—provide a template for studying other classes of cryptographic functions.
Acknowledgments
The third-named author (PS) thanks the first-named author (DB) for the invitation to Università degli Studi di Perugia and the excellent working conditions. The research of the first-named author (DB) was supported by the Italian National Group for Algebraic and Geometric Structures and their Applications (INdAM - GNSAGA Project, CUP E53C24001950001).
Appendix A Polynomial definitions
We list below the functions $f(\alpha,\beta,X,Y,Z)$, $g(\beta,X,Y,Z)$, $h(\beta,X,Y,Z)$, $a_{i}(\beta,Y,Z)$, $i=0,\ldots,21$, needed in Proposition 6.1: